OpenStudy (anonymous):

The region between the graphs of y=x^2 and y=5x is rotated around the line y=25. Find the volume

OpenStudy (anonymous):

so far i have integral 0 to 25 ((5-sqrt(x)^2)-(5-(x/5))^2

OpenStudy (mathmate):

The axis is y=25, and I don't see you subtracting anything from 25. You can also opt to use the washer method, if that's what you're doing at school.

OpenStudy (anonymous):

so how do i solve this can u do it step by step?

OpenStudy (mathmate):

Do you prefer the shell method or washer method?

OpenStudy (mathmate):

Have you done either of the two methods in your course?

OpenStudy (anonymous):

washer

OpenStudy (mathmate):

There you go, then the integral would be the difference of two squares. The first one is the "outer radius", meaning from the axis to the outmost part of the area x^2 in our case. This is (25-x^2)^2. The second square is from the axis to the edge of the "hole", i.e. empty space we don't want to integrate, in our case this is (25-5x)^2. Multiply all by pi to get: \[\pi \int\limits_{0}^{5}((25-x^2)^2-(25-5*x)^2)dx\] Does that look familiar?

OpenStudy (anonymous):

how do u get 25-5x

OpenStudy (mathmate):

Notice that we are integrating along the x-direction, so dx would represent a vertical slice. (25-5x) is the distance between the axis of rotation and the top of our region. This represents the "hole" of the washer, or the area that we need to subtract from a disk (to make a washer.

OpenStudy (anonymous):

OOOHHH

OpenStudy (anonymous):

yes ok

OpenStudy (anonymous):

Find the volume of the solid obtained by rotating the region bounded by the given curves about the line x=−4 y=x^2, x=y^2

OpenStudy (anonymous):

how about something like that using shell, now this i have no idea how to set up using shell but i was advised to

OpenStudy (anonymous):

?

OpenStudy (mathmate):

Sorry, I'm back!

OpenStudy (anonymous):

Find the volume of the solid obtained by rotating the region bounded by the given curves about the line x=−4 y=x^2, x=y^2

OpenStudy (anonymous):

how about something like that using shell, now this i have no idea how to set up using shell but i was advised to

OpenStudy (mathmate):

This one you can do either shell or washer.

OpenStudy (mathmate):

You could start with making a plot to find the points of intersection.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

give me a sec

OpenStudy (anonymous):

0 to 1?

OpenStudy (mathmate):

Yep! I suggest you do this by the method of washers to practice what we just did. If you succeed, then we can use the shell method as a check.

OpenStudy (anonymous):

ok lets do it

OpenStudy (mathmate):

Can you figure out the outer and inner radius?

OpenStudy (anonymous):

tha was wat i was just about to ask

OpenStudy (anonymous):

y=x^2 is outter

OpenStudy (mathmate):

Even before that, can you tell me which functions you used to find the intersection?

OpenStudy (mathmate):

There is x^2, is the other one sqrt(x)?

OpenStudy (mathmate):

Sqrt(x) is the part of x=y^2 above the x-axis. We like to do that because x=y^2 is not strictly a function.

OpenStudy (mathmate):

Are you still there?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok i understand

OpenStudy (mathmate):

So the axis is at y=-4, right?

OpenStudy (anonymous):

yes

OpenStudy (mathmate):

Can you give me the "outer" radius, meaning the expression that gives the distance of the top part of the area relative to the axis (y=-4, which is below the area).

OpenStudy (anonymous):

i think its stil y=x^2

OpenStudy (mathmate):

x^2 is the lower curve. From the axis (y=-4), sqrt(x) is the outer (upper) curve. Do you think so?

OpenStudy (anonymous):

1 second

OpenStudy (anonymous):

yes i saw it once i zoomed in

OpenStudy (anonymous):

sqrt

OpenStudy (anonymous):

of x

OpenStudy (mathmate):

So the distance for the "outer" radius is therefore sqrt(x)+4, right?

OpenStudy (anonymous):

yes

OpenStudy (mathmate):

And the "inner" radius (referrng to the lower curve x^2) is then....

OpenStudy (anonymous):

so (sqrt(x)-4)-x^2

OpenStudy (mathmate):

Not really, for the inner, we're just interested in the x^2 curve, so it would be x^2+4

OpenStudy (anonymous):

(sqrt(x)+4)-(x^2+4)

OpenStudy (mathmate):

Almost there, you need to square them, and multiply everything by pi. This way, pi(outer)^2 is the big disk, and pi(inner)^2 is the small (empty) disk. The difference becomes a washer that you want to calculate.

OpenStudy (anonymous):

yes ok, let me put it into my program to see wat i get, give me a minute

OpenStudy (anonymous):

Fantastic it was right

OpenStudy (anonymous):

i have one more but i think i know wat the upper an lower radius' are

OpenStudy (mathmate):

Give it a try!

OpenStudy (anonymous):

Find the volume of the solid obtained by rotating the region bounded by the curves y=x^4, y=1

OpenStudy (mathmate):

about ?

OpenStudy (anonymous):

outter radius is 4-x^2?

OpenStudy (anonymous):

inner is 4-1?

OpenStudy (mathmate):

Is it about y=+4?

OpenStudy (anonymous):

yes

OpenStudy (mathmate):

x^2 or x^4?

OpenStudy (anonymous):

Find the volume of the solid obtained by rotating the region bounded by the curves y=x^4, y=1 about the line y=4

OpenStudy (mathmate):

What you suggested seem OK to me, just change the x^2 to x^4, must have been a typo.

OpenStudy (anonymous):

wait why x^4?

OpenStudy (mathmate):

I thought your question asked for x^4, right?

OpenStudy (anonymous):

oh yes lol

OpenStudy (anonymous):

my bad

OpenStudy (mathmate):

no problem! lol

OpenStudy (anonymous):

i got it wrong though

OpenStudy (anonymous):

i got 248/45pi

OpenStudy (mathmate):

Can you show me what you put in?

OpenStudy (anonymous):

\[\int\limits_{0}^{1}(4-x^4)^2-(4-1)^2)\]

OpenStudy (mathmate):

Let met see...

OpenStudy (mathmate):

It's just the pi missing, unless the answer does not correspond to the question.

OpenStudy (mathmate):

Let's review the question: Ah, what are your limits?

OpenStudy (mathmate):

x^4 is an even function!

OpenStudy (anonymous):

0 to 1

OpenStudy (mathmate):

I suppose you have not yet graphed the function to find the enclosed area?

OpenStudy (mathmate):

An even function typically take - to +, like -1 to +1.

OpenStudy (anonymous):

ok so what did u get as the interval?

OpenStudy (mathmate):

-1 to +1, that's where the two functions intersect.

OpenStudy (anonymous):

yes u were right

OpenStudy (mathmate):

So everything's good?

OpenStudy (anonymous):

yes thanks a lot

OpenStudy (mathmate):

Let's resume: 1. find the shape of the area to rotate, and find the limits (graphing). 2. if using the washer method, find the outer and inner radii. 3. integrate using the correct limits, and don't forget pi. OK?

OpenStudy (anonymous):

hahaha but i finished my homework professor

OpenStudy (mathmate):

That's OK, you'll still have exams to come, that's when YOUR professor has the last laugh! lol

OpenStudy (mathmate):

Anyway, have a great weekend, and enjoy!

OpenStudy (anonymous):

thanks a lot

OpenStudy (mathmate):

You're welcome!

OpenStudy (anonymous):

when are u usually on?

OpenStudy (mathmate):

Evenings.

OpenStudy (anonymous):

ok catch u later

OpenStudy (mathmate):

Yep!